# Chain-of-Instructions: Compositional Instruction Tuning on Large Language Models

### v2

@@ Abstract
Fine-tuning large language models (LLMs) with a collection of large and diverse instructions has improved the model’s generalization to different tasks, even for unseen tasks. However, most existing instruction datasets include only single instructions, and they struggle to follow complex instructions composed of multiple subtasks. In this work, we propose a novel concept of compositional instructions called chain-of-instructions (CoI), where the output of one instruction becomes an input for the next like a chain. Unlike the conventional practice of solving single instruction tasks, our proposed method encourages a model to solve each subtask step by step until the final answer is reached. CoI-tuning (i.e., fine-tuning with CoI instructions) improves the model’s ability to handle instructions composed of multiple subtasks as well as unseen composite tasks such as multilingual summarization. Overall, our study find that simple CoI tuning of existing instruction data can provide consistent generalization to solve more complex, unseen, and longer chains of instructions. [^1] We will publicly release our code and CoI data.
Chain-of-Instructions: Compositional Instruction Tuning on
Large Language Models
Shirley Anugrah Hayati thanks:   Work was partially done during internship at Amazon.  [Uncaptioned image]      Taehee Jung[Uncaptioned image]      Tristan Bodding-Long[Uncaptioned image] Sudipta Kar[Uncaptioned image]      Abhinav Sethythanks:   Work was done while at Amazon.[Uncaptioned image]      Joo-Kyung Kim[Uncaptioned image]      Dongyeop Kang[Uncaptioned image] [Uncaptioned image]University of Minnesota      [Uncaptioned image]Amazon AGI      [Uncaptioned image]Grammarly {hayat023, dongyeop}@umn.edu {jungtaeh, boddint, sudipkar, sethya, jookyk}@amazon.com abhinav.sethy@grammarly.com
1 Introduction
@@
Large language models (LLMs) have demonstrated impressive performance in various tasks, from conventional NLP downstream tasks, such as machine translation and summarization, to open-ended tasks, such as writing an outline for blog posts and giving tips for presentation, when fine-tuned on human-like instructions Ouyang et al. (2022); Wang et al. (2022); Conover et al. (2023); Mishra et al. (2022). These models excel at single instruction tasks, however, their ability to handle complex and compositional instructions is less explored.
@@
Figure 1: Example Chain-of-Instruction (CoI)s in our dataset. In this example, the output of a summarization subtask can become the input of a title generation subtask, and the output of the title generation task can be an input for a style transfer or translation task. The thickness of arrows denotes the probability of whether we can compose two corresponding subtask instructions. X denotes that we cannot compose these subtask instructions due to format mismatch. $I_{k}$ is $k^{th}$ instruction and $O_{k}$ is $k^{th}$ output.
@@
A compositional instruction contains a series of sequential subtasks as a subtask’s output becomes the input for the next one in a chained manner as shown in Figure 1. We call this problem as Chain-of-Instructions or shortly CoI. What subtasks can be composed more naturally than others?
@@
In Figure 1, we can see that some tasks can be composed together, such as inputting a summary to a title generation task, while some tasks cannot be composed together, e.g., a summary as an input for a Data-to-Text task. Some tasks have higher probability of being able to be composed, such as generating a title from a summary compared to converting numbers in a summary since sometimes a summary does not contain a number. Figure 2 illustrates a more detailed example of CoI. The given instruction “Generate a blog-like title in French” can be decomposed into three chained sub-instructions:
@@
Self-Instruct Wang et al. (2023) 52k Daily QA-like tasks
Dolly Conover et al. (2023) 15k Daily QA-like tasks
Super-NaturalInstruct Wang et al. (2022) 1.6k NLP tasks
Faith and Fate Dziri et al. (2023) N/A Math, logic, programming
Compositional Semantic Drozdov et al. (2022) N/A CFQ, COGS, Parsing
MuSiQue Trivedi et al. (2022) 24.8k Multi-hop QA
Table 1: A comparison of our work with existing related works. As some previous works do not contribute a new dataset, the dataset size is shown as N/A. For instruction datasets, data size refers to the number of instructions, not task instances (input-output pair). More prior work is studied in §5.
@@
Our study investigates whether LLMs can handle compositional instructions effectively and whether models tuned with compositional instructions can be generalized to solve more complex, unseen, or longer chains of instructions. We first create a new CoI dataset with our proposed LLM-based compositionality checker, and then evaluate our model’s performance in handling (1) traditional single instructions and (2) compositional instructions.
Our work is closely related to other instruction-tuning works and compositional studies in NLP, as summarized in Table 1. Wang et al. (2023); Conover et al. (2023); Wang et al. (2022) propose new instruction datasets, but they only handle single instruction problems. Although our approach draws inspiration from Chain-of-Thought (CoT) prompting Wei et al. (2022b) or Least-to-Most prompting Zhou et al. (2022), our CoI is not a prompting technique but a collection of chained instructions validated by an LLM, showing generalization in solving complex and compositional problems. Our contributions are as follows:
@@
2 Chain-of-Instructions
2.1 Formulation
@@
Two single Instructions from Sup-NatIns Wang et al. (2022)
Task category: Title Generation Style Transfer
Original instruction:
“You are given a paragraph about
various topics. Your task is to generate
a title for the given paragraph.”
“Convert the style of the given text to be
suitable for a blog. Please consider
the following points.. ”
Summarized instruction: “Generate a title” “Convert the text into a blog-style text”
Input:
Ratatouille is a traditional stew made with
summer vegetables. When cooking Ratatouille..
You can put an arrow on an elbow connector
if you use the app.
Output: Ratatouille: A Taste of Summer’s farewell
Hey, there! Ever wondered if you can spice
up your elbow connectors with a sleek arrow?
Compositional Instruction: Title Generation $\rightarrow$ Style Transfer
Instruction: “Generate a title and then convert the title into a blog-style title”
Input: Ratatouille is a traditional stew made with summer vegetables. When cooking Ratatouille…
Output: Ratatouille: A Taste of Summer’s Farewell
Delight in Ratatouille: Savor the Flavors of a Summer Farewell
Table 2: Single instruction examples (top) and compositional instruction example in with CoI variants (bottom). The first subtask is to “generate a title” and the second subtask is to “convert the text into a blog-style text.” Red text refers to the first subtask, and green text for the second subtask. The input for the first task in the compositional instruction is the same, which is a text about Ratatouille.
@@
An example of CoI is shown in Table 2.
2.2 Automatic Dataset Creation Pipeline
@@
We curate a new compositional instruction dataset from existing single task instruction dataset: Super-NaturalInstructions (Sup-NatIns) Wang et al. (2022). We select Sup-NatIns as the seed dataset because it contains a wide variety of tasks (1,616 unique tasks) from 76 categories, including text categorization, summarization, and machine translation. Each category contains many different NLP tasks. For example, under the text categorization category, there exist sarcasm detection and politeness classification tasks. Each task in Sup-NatIns contains human-written descriptions that can be considered as instructions and instances as pairs of inputs and outputs. We only select tasks with English (1,341 unique tasks) as their input language to make sure that the chain is connected. For our single-task instruction tuning data (CoI1), we randomly sample 10 input-output pairs, resulting in 13,410 instances.
Figure 3: Data creation process for CoI2. We use an LLM for both instruction summarization and composability check. The right column shows an example instance of our chain-of-instruction dataset. Output 1 in Step 2 comes from the original SupNatInst data.
@@
Composing two single instructions poses a challenge due to their lengthy and specific descriptions, and differing output formats. Figure 3 illustrates a two-step process for creating a compositional instruction dataset with the help of an LLM as elaborated in the next paragraphs. Here we use GPT 3.5 Turbo Ouyang et al. (2022). [^2] We use GPT 3.5 because of its reasonable price and at the time of we examine that the quality of the result is good enough. However, this data creation procedure can be reproducible with other strong LLMs as well.
@@
The task instructions in Sup-NatIns are lengthy and detailed, which may deviate from real human-like instructions. With the same dataset (Sup-NatIns), Yin et al. (2023) find that 60% of tokens can be removed with comparable, if not better, model performance. Thus, we use the LLM to shorten each instruction in the Sup-NatIns dataset. This step reduces the average number of words in the Sup-NatIns descriptions from 62.19 to 14.33.
@@
To generate compositional instructions from single instructions, we undertake a two-step process: (1) validity check and (2) generate the output for the second (or third) subtask. Validity check is done to examine whether two subtasks are composable. We first filter out non-composable tasks with heuristics developed by the authors’ knowledge (A.7). For example, classification tasks can only be the last subtask when composing a pair of tasks. After applying these heuristics, we additionally check whether LLM can generate the output for the second instruction based on the first instruction’s input. If yes, we treat the pair as composable. [^3] Prompt for this step is available in A.2.2.
@@
To examine the quality of LLM’s composability check, we randomly sampled 100 instances and manually inspected which composed instructions are valid. We find that 75% are valid composed instructions. For CoI3, similarly we randomly sampled 100 instances and found that 59% are valid compositions. [^4] Such error rates are often found in LLM-generated data Das et al. (2024); Wang et al. (2023).
2.3 CoI Dataset
Table 3 shows the data statistics of CoI datasets. In chain length 2, we obtain 970 unique category pairs; in chain length 3, we obtain 418 unique category triplets. In each pair or triplet, we randomly select at most three instances and divide them into training and testing sets. For the longer chains (4, 5), we only use them for testing. Please find Appendix A.1 for the detailed statistics.
Figure 4 shows a t-SNE plot when we embed subtask instructions of frequent CoI2 instructions using SentenceBERT Reimers and Gurevych (2019) with DistilRoberta Sanh et al. (2019). [^5] We only select instruction pairs that appear more than 7 times, and 9 is a max number of occurrences in CoI2 dataset. We find generation tasks such as paraphrasing and question generation can be compiled as both the first and second subtasks, except for problems involving specific input formats, such as code to text or data to text, which can only be compiled as the second subtask. On the other hand, close-ended problems (e.g., POS tagging or grammar error detection) mostly appear as the second subtask.
@@ 3 Experiment Setup
We fine-tune the base models of Alpaca-7B Taori et al. (2023) and Mistral-7B-Instruct Jiang et al. (2023). Since both models are open-sourced single instruction-tuned models which are widely used, they are suitable to be compared with CoI-tuned models.
@@
  • Off-the-shelf version of Alpaca-7B Taori et al. (2023) model and Mistral-7B-Instruct model without fine-tuning (Base).
  • The same non-finetuned Alpaca and Mistral with chain-of-thought prompting Wei et al. (2022b) (CoT) with seven-shot demonstrations and least-to-most prompting Zhou et al. (2022) (LtM).
@@
Table 4: Rouge-L results on compositional instruction test sets and BIG-Bench Hard (BBH). Base refers to the non-fine-tuned base models, CoT = chain-of-thought prompting on base models, LtM = least-to-most prompting on base models. The best scores are bold and the second best scores are underlined.
Metrics
For our evaluation metric, we report Rouge-L Lin (2004), following Wang et al. (2022) since it can be used for measuring different text generation tasks. Note that automatic metrics, such as Rouge score focus primarily on surface-level overlap and it may not capture the correctness and preferred nuance of the output. Thus, we also have human evaluation to do blind pairwise comparison between the outputs from the baseline and from our best CoI models.
Test sets
  • BIG-Bench Hard For the single instruction test set, we use BIG-Bench Hard Suzgun et al. (2022), a collection of 27 challenging tasks, such as date understanding and evaluating the truth value of Boolean expressions (section A.4), and each task has $\leq$ 250 instances. BIG-Bench Hard subset enables us to evaluate the model’s performance on diverse and challenging NLP tasks with clear single instructions and associated input-output pairs.
  • Downstream Task In addition to CoI test sets, we examine the usefulness of CoI on the downstream task of multilingual summarization using WikiLingua Ladhak et al. (2020) [^7] This is a multilingual dataset based on WikiHow [^6] https://www.wikihow.com for abstractive summarization in 18 languages (section A.5). WikiHow articles provide step-by-step instructions to complete procedural tasks on different topics, and each step includes a one-sentence summary as well.. In our experiment, we select source-target language pairs ; English-to-French (WikiLingua-en-fr) and Spanish-to-English (WikiLingua-es-en) and randomly sample 300 test instances for each. Given an input content from source language $L_{src}$, we aim to generate a summary in target language $L_{tgt}$. This task is similar to a 2-instruction problem as we summarize first and then translate. Note that CoI training data only contains translation tasks from English to Punjabi, German, and Catalan, thus, selected source-target pairs are unseen in CoI training set.
Mistral Alpaca
Base CoI Base CoI
Test Set: CoI2
Subtask 1 1.32 90.50 13.98 84.16
Subtask 2 2.40 49.21 7.02 45.57
Test Set: CoI3
Subtask 1 18.04 81.49 9.56 91.77
Subtask 2 6.82 68.65 2.13 71.67
Subtask 3 6.93 32.73 3.30 35.52
Table 5: Rouge-L results on intermediate tasks. CoI models refer to best models of CoI: CoI12 model if the test set=CoI2 and CoI123 model if the test set=CoI3. The best scores are marked as bold.
4 Results
We conduct experiments to measure the performance of CoI-tuned models on our compositional instructions (§4.1), and the generalization capability to difficult single instructions (§4.2), and longer-chain instructions $\sigma=\{4,5\}$4.3), and the application to an existing downstream task (§4.4). [^8] We also conduct an ablation study to see if the correctness of second and third subtask outputs matter in the Table 10 in Appendix. We see degrading performance of models fine-tuned with incorrect outputs, showing the importance to have the correct output for the subtasks during training.
@@ 4 Results
4.1 Performance on In-domain Composite Tasks (CoI2,3)
@@
As we evaluate our CoI models’ performance against the baselines on multi-CoI test sets, we find that both Mistral and Alpaca fine-tuned on CoI12 instructions perform the best for CoI2-test (Table 4)  [^9] Fine-tuning details in Appendix A.3.. Similarly, for CoI3-test, both CoI123 Mistral and Alpaca perform the best. All models fine-tuned on compositional instructions generally outperform the baselines, except for CoI3-tuned Alpaca. This model performs slightly worse than the CoI1-tuned Alpaca on CoI2 test set. We hypothesize that this happens because instructions in CoI3 become very long, thereby it becomes harder for the model to generalize without CoI2 and CoI1 examples. As a result, models only fine-tuned on CoI3 tend to generate long sentences with hallucinations as in Table 6.
@@
We examine how CoI models perform for each subtask in the compositional instruction. To do this, we compare the results from the best version of CoI (CoI12 for CoI2 test set, CoI123 for CoI3 test set) against the non-finetuned baseline models. Since there is no clear boundary to distinguish the first subtask output and the second subtask output in the baseline’s outputs, we use an LLM to separate the responses. Given the subtask instruction and the output, we ask the LLM to decide which span of the output text responds to the subtask instruction. To remove the possibility of LLM’s hallucination being counted as part of the output, we only include LLM’s output span when it appears in the baseline’s output. When LLM deems that the output is incorrect, we assign Rouge = 0 because we do not know whether this output refers to the first subtask or second subtask.
Table 5 and Table 12 show results of CoI models and baseline on CoI2 and CoI3 test sets. In general, CoI models outperform the baseline for both test sets, with the best results on the first subtask and then followed by the second subtask and the third subtask. However, it is interesting to see that CoI’s performance on the second subtask for CoI3 test set is higher than CoI’s performance on the second subtask for CoI2. We conjecture that it happens because the tasks that are composable as a previous subtask can be easier to solve or knowing the subsequent subtask helps improving the performance.
@@
4.2 Generalization to Unseen Single Tasks
To assess whether adding compositional instructions help improve the model’s performance on unseen and difficult single instruction tasks, we tested CoI-tuned models on BIG-Bench Hard (BBH). CoI123-tuned Mistral performs the best (Rouge: 29.39) as shown in Table 4. For Alpaca, the model fine-tuned on CoI12 is also better than the baseline and achieves a higher Rouge score of 28.80. This confirms that having compositional instructions helps the model to understand hard single instruction problems as well.
4.3 Generalization to Longer Chains (CoI4,5)
In this experiment, we examine whether our CoI models can generalize to longer chains. We run inference on CoI4 and CoI5 test sets using CoI1, CoI12, and CoI123-tuned Alpaca. [^10] Mistral results are in Table 11 in Appendix A.9. As shown in Figure 6, longer chains ($\sigma=2,3$) in the training set help the model to understand unseen longer chain ($\sigma=4,5$) in the test set as well. Moreover, the performance does not drop as high as CoI1 or the baseline non-fine-tuned models that do not learn the chaining reasoning. We posit that the knowledge of compositional instructions in the training set, even though the length of the chain is shorter than 4 or 5, still helps the model to understand the composed tasks.
@@ 4.3 Generalization to Longer Chains (CoI4,5)
4.4 Generalization to Downstream Composite Tasks
@@ 5 Related Work
There has been a notable surge in research focused on fine-tuning LLMs using human instructions. Efrat and Levy (2020) examined LLMs’ ability to follow natural language instructions compared to crowd-workers. Wei et al. (2022a); Sanh et al. (2021) have transformed NLP task descriptions into human-like language instructions and showed that LLMs fine-tuned with those instructions have generalizable capability toward unseen tasks Chung et al. (2024). Subsequently, many studies have emerged to create new instruction datasets aimed at training models in instruction-tuning paradigm: some instruction datasets are fully written by humans Wang et al. (2022); Conover et al. (2023), the others are written with the help of LLMs Honovich et al. (2023); Wang et al. (2023); Taori et al. (2023); some instructions are NLP-specific Mishra et al. (2022); Wang et al. (2022); Weller et al. (2020), and the others are designed to respond to general-purpose instructions Ouyang et al. (2022); Wang et al. (2023). Raheja et al. (2023) hypothesizes that task-specific instructions enforce the model generalization to unseen tasks. These prior studies only work on single instruction datasets, so we construct a new compositional dataset upon Wang et al. (2022)’s Super-NaturalInstruction. Our work is also related to several past works which have leveraged LLMs to generate training data Schick and Schütze (2021); Yu et al. (2023), and some of them specifically use LLMs for generating instruction data Sun et al. (2023); Peng et al. (2023); Shao et al. (2023). Nevertheless, our CoI data generation framework differs from previous works as we use LLMs to determine the composability of individual instructions, and then generate responses for subsequent subtask instructions.

### v3

@@ Abstract
Fine-tuning large language models (LLMs) with a collection of large and diverse instructions has improved the model’s generalization to different tasks, even for unseen tasks. However, most existing instruction datasets include only single instructions, and they struggle to follow complex instructions composed of multiple subtasks. In this work, we propose a novel concept of compositional instructions called chain-of-instructions (CoI), where the output of one instruction becomes an input for the next like a chain. Unlike the conventional practice of solving single instruction tasks, our proposed method encourages a model to solve each subtask step by step until the final answer is reached. CoI-tuning (i.e., fine-tuning with CoI instructions) improves the model’s ability to handle instructions composed of multiple subtasks as well as unseen composite tasks such as multilingual summarization. Overall, our study finds that simple CoI tuning of existing instruction data can provide consistent generalization to solve more complex, unseen, and longer chains of instructions. Our code and data are available at https://github.com/amazon-science/chain-of-instructions.
Introduction
@@
Large language models (LLMs) have demonstrated impressive performance in various tasks, from conventional NLP downstream tasks, such as machine translation and summarization, to open-ended tasks, such as writing an outline for blog posts and giving tips for presentation, when fine-tuned on human-like instructions (Ouyang et al. 2022; Wang et al. 2022; Conover et al. 2023; Mishra et al. 2022). These models excel at single instruction tasks, but their ability to handle complex and compositional instructions is less explored.
@@
Figure 1: Chain-of-Instructions (CoI) example. The summarization output can be an input for a title generation subtask; the output of the title generation can be an input for style transfer or translation subtasks. Arrow thickness denotes the probability of instruction composability. X means that these subtasks cannot be composed due to format mismatch. $I_{k}$ is $k^{th}$ instruction and $O_{k}$ is $k^{th}$ output.
@@
A compositional instruction contains a series of sequential subtasks, as the output of one subtask becomes the input of the next one in a chained manner as shown in Figure 1. We call this problem as Chain-of-Instructions or shortly CoI. We examine what subtasks can be composed more naturally than others.
@@
In Figure 1, we can see that some tasks can be composed together, such as input a summary to a title generation task, while some tasks cannot be composed together, e.g., a summary as input for a Data-to-Text task. Some tasks have a higher probability of being able to be composed, such as generating a title from a summary compared to converting numbers in a summary since sometimes a summary does not contain a number. Figure 2 illustrates a more detailed example of CoI. The given instruction “Generate a blog-like title in French” can be decomposed into three chained sub-instructions:
@@
Self-Instruct (Wang et al. 2023) 52k Daily QA-like tasks
Dolly (Conover et al. 2023) 15k Daily QA-like tasks
Super-NaturalInstruct (Wang et al. 2022) 1.6k NLP tasks
Faith and Fate (Dziri et al. 2023) N/A Math, logic, programming
Compositional Semantic (Drozdov et al. 2022) N/A CFQ, COGS, Parsing
MuSiQue (Trivedi et al. 2022) 24.8k Multi-hop QA
Table 1: A comparison of our work with existing related works. As some previous works do not contribute a new dataset, the dataset size is shown as N/A. For instruction datasets, data size refers to the number of instructions, not task instances (input-output pair). More prior work is studied in §Related Work.
@@
When these sub-instructions are composed, we call it a compositional instruction or chain-of-instructions. Our study investigates whether LLMs can handle compositional instructions effectively and whether models tuned with compositional instructions can be generalized to solve more complex, unseen, or longer chains of instructions. We first create a new CoI dataset with our proposed LLM-based compositionality checker, and then evaluate our model’s performance in handling (1) traditional single instructions and (2) compositional instructions.
Our work is closely related to other instruction-tuning works and compositional studies in NLP, as summarized in Table 1. Wang et al. (2023); Conover et al. (2023); Wang et al. (2022) propose new instruction datasets, but they only handle single instruction problems. Although our approach draws inspiration from Chain-of-Thought (CoT) prompting (Wei et al. 2022b) or Least-to-Most prompting (Zhou et al. 2022), our CoI is not a prompting technique but a collection of chained instructions validated by an LLM, showing generalization in solving complex and compositional problems. Our contributions are as follows:
@@
Chain-of-Instructions
Formulation
@@
Automatic Dataset Creation Pipeline
@@
We curate a new compositional instruction dataset from existing single task instruction dataset: Super-NaturalInstructions (Sup-NatIns) (Wang et al. 2022). We select Sup-NatIns as the seed dataset because it contains a wide variety of tasks (1,616 unique tasks) from 76 categories, including text categorization, summarization, and machine translation. Each category contains many different NLP tasks. For example, under the text categorization category, there exist sarcasm detection and politeness classification tasks. Each task in Sup-NatIns contains human-written descriptions that can be considered as instructions and instances as pairs of inputs and outputs. We only select tasks with English (1,341 unique tasks) as their input language to make sure that the chain is connected. For our single-task instruction tuning data (CoI1), we randomly sample 10 input-output pairs, resulting in 13,410 instances.
Figure 3: Data creation for CoI2. We use an LLM for both instruction summarization and composability check. The right column shows an example instance of our chain-of-instruction dataset. Output 1 in Step 2 comes from the original SupNatInst data.
@@
Composing two single instructions poses a challenge due to their lengthy and specific descriptions, and differing output formats. Figure 3 illustrates a two-step process for creating a compositional instruction dataset with the help of an LLM as elaborated in the following paragraphs. Here we use GPT 3.5 Turbo (Ouyang et al. 2022) because of its reasonable price and at the time, we examine that the quality of the result is good enough. However, this data creation procedure can be reproducible with other strong LLMs as well.
@@
The task instructions in Sup-NatIns are lengthy and detailed, which may deviate from real human-like instructions. With the same dataset (Sup-NatIns), Yin et al. (2023) find that 60% tokens can be removed with comparable, if not better, model performance. Thus, we use the LLM to shorten each instruction in the Sup-NatIns dataset. This step reduces the average number of words in the Sup-NatIns descriptions from 62.19 to 14.33.
@@
To generate compositional instructions from single instructions, we perform a two-step process: (1) validity check and (2) generate the output for the second (or third) subtask. The validity check is performed to examine whether two subtasks are composable. We first filter out non-composable tasks with heuristics developed by the authors’ knowledge (the Heuristics for Validity Check section in the Appendix). For example, classification tasks can only be the last subtask when composing a pair of tasks. After applying these heuristics, we additionally check whether LLM can generate the output for the second instruction based on the input of the first instruction. If so, we treat the pair as composable. [^1] Prompt for this step is available in the Appendix.
@@
To examine the quality of LLM’s composability check, we randomly sampled 100 instances and manually inspected which composed instructions are valid. We find that 75% are valid composed instructions. For CoI3, similarly we randomly sampled 100 instances and found that 59% are valid compositions. Such error rates are often found in LLM-generated data (Das et al. 2024; Wang et al. 2023).
CoI Dataset
Table 2 shows the data statistics of CoI datasets. In chain length 2, we obtain 970 unique category pairs; in chain length 3, we obtain 418 unique category triplets. In each pair or triplet, we randomly select at most three instances and divide them into training and testing sets. For the longer chains (4, 5), we only use them for testing. Please find Appendix CoI Dataset Statistics for the detailed statistics.
Figure 4 shows a t-SNE plot when we embed subtask instructions of frequent CoI2 instructions using SentenceBERT (Reimers and Gurevych 2019) with DistilRoberta (Sanh et al. 2019). [^2] We only select instruction pairs that appear more than 7 times, and 9 is a max number of occurrences in CoI2 dataset. We find generation tasks such as paraphrasing and question generation can be compiled as both the first and second subtasks, except for problems involving specific input formats, such as code to text or data to text, which can only be compiled as the second subtask. On the other hand, close-ended problems (e.g., POS tagging or grammar error detection) mostly appear as the second subtask.
@@
We fine-tune the base models of Alpaca-7B (Taori et al. 2023) and Mistral-7B-Instruct (Jiang et al. 2023). Since both models are open-sourced single instruction-tuned models which are widely used, they are suitable to be compared with CoI-tuned models.
@@
  • Off-the-shelf version of Alpaca-7B (Taori et al. 2023) model and Mistral-7B-Instruct model without fine-tuning (Base).
  • The same non-finetuned Alpaca and Mistral with chain-of-thought prompting (Wei et al. 2022b) (CoT) with seven-shot demonstrations and least-to-most prompting (Zhou et al. 2022) (LtM).
@@
Metrics
For our evaluation metric, we report Rouge-L (Lin 2004), following Wang et al. (2022) and LLM (gpt-4o-mini) as a preference judge. Rouge can be used to assess various text generation tasks and using LLM as a judge has been widely adopted in NLP research (Liu et al. 2023; Fu et al. 2024). We also have human evaluation to perform blind pairwise comparison between the outputs from the baseline and from our best CoI models.
Test sets
  • BIG-Bench Hard For the single instruction test set, we use BIG-Bench Hard (Suzgun et al. 2022), a collection of 27 challenging tasks, such as date understanding and evaluating the truth value of Boolean expressions, and each task has $\leq$ 250 instances. BIG-Bench Hard subset enables us to evaluate the model’s performance on diverse and challenging NLP tasks with clear single instructions and associated input-output pairs.
  • Downstream Task In addition to CoI test sets, we examine the usefulness of CoI on the downstream task of multilingual summarization using WikiLingua (Ladhak et al. 2020), which is a multilingual dataset based on WikiHow [^3] https://www.wikihow.com for abstractive summarization in 18 languages. WikiHow articles provide step-by-step instructions to complete procedural tasks on different topics, and each step includes a one-sentence summary as well. In our experiment, we select source-target language pairs ; English-to-French (WikiLingua-en-fr) and Spanish-to-English (WikiLingua-es-en) and randomly sample 300 test instances for each. Given an input content from source language $L_{src}$, we aim to generate a summary in target language $L_{tgt}$. This task is similar to a 2-instruction problem as we summarize first and then translate. Note that CoI training data only contains translation tasks from English to Punjabi, German, and Catalan, thus, selected source-target pairs are unseen in CoI training set.
Mistral Alpaca
Base CoI Base CoI
Test Set: CoI2
Subtask 1 1.32 90.50 13.98 84.16
Subtask 2 2.40 49.21 7.02 45.57
Test Set: CoI3
Subtask 1 18.04 81.49 9.56 91.77
Subtask 2 6.82 68.65 2.13 71.67
Subtask 3 6.93 32.73 3.30 35.52
Table 3: Rouge-L results on intermediate tasks. CoI models refer to best models of CoI: CoI12 model if the test set=CoI2 and CoI123 model if the test set=CoI3.
Results
We conduct experiments to measure the performance of CoI-tuned models on our compositional instructions (§Performance on In-domain Composite Tasks (CoI2,3)), and the generalization capability to difficult single instructions (§Generalization to Unseen Single Tasks), and longer-chain instructions $\sigma=\{4,5\}$Generalization to Longer Chains (CoI4,5)), and the application to an existing downstream task (§Generalization to Downstream Composite Tasks). We also conduct an ablation study to see if the correctness of second and third subtask outputs matter in the Table 9 in Appendix. We see degrading performance of models fine-tuned with incorrect outputs, showing the importance to have the correct output for the subtasks during training.
@@ Results
Table 4: Rouge-L results on compositional instruction test sets and BIG-Bench Hard (BBH). Base refers to the non-fine-tuned base models, CoT = chain-of-thought prompting on base models, LtM = least-to-most prompting on base models. The best scores are marked as bold.
@@
Performance on In-domain Composite Tasks (CoI2,3)
@@
As we evaluate our CoI models’ performance against the baselines on multi-CoI test sets, we find that both Mistral and Alpaca fine-tuned on CoI12 instructions perform the best for CoI2-test (Table 4)  [^4] Fine-tuning details in Appendix.. Similarly, for CoI3-test, both CoI123 Mistral and Alpaca perform the best. All models fine-tuned on compositional instructions generally outperform the baselines, except for CoI3-tuned Alpaca. This model performs slightly worse than the CoI1-tuned Alpaca on CoI2 test set. We hypothesize that this happens because instructions in CoI3 become very long, thereby it becomes harder for the model to generalize without CoI2 and CoI1 examples. As a result, models only fine-tuned on CoI3 tend to generate long sentences with hallucinations as in Table 5.
In the LLM-as-a-judge experiment, we evaluate the performance of the best CoI models on CoI2-test and CoI3-test against the best baseline, CoI1. On CoI2-test, the LLM prefers 69.90% of Alpaca CoI12’s outputs and 70.92% of Mistral CoI12’s outputs over the baseline. Similarly, on CoI3-test, the LLM favors Alpaca CoI123’s outputs and Mistral CoI123’s outputs over the baseline by 81.04% and 60.00%, respectively.
@@
We examine how CoI models perform for each subtask in the compositional instruction. To do this, we compare the results from the best version of CoI (CoI12 for CoI2 test set, CoI123 for CoI3 test set) against the non-finetuned baseline models. Since there is no clear boundary to distinguish the first subtask output and the second subtask output in the baseline’s outputs, we use an LLM to separate the responses. Given the subtask instruction and the output, we ask the LLM to decide which span of the output text responds to the subtask instruction. To remove the possibility of LLM’s hallucination being counted as part of the output, we only include LLM’s output span when it appears in the baseline’s output. When LLM deems that the output is incorrect, we assign Rouge = 0 because this output could refer to the first subtask or second subtask.
Table 3 and Table 11 show results of CoI models and baseline on CoI2 and CoI3 test sets. In general, CoI models outperform the baseline for both test sets, with the best results on the first subtask and then followed by the second subtask and the third subtask. However, it is interesting to see that CoI’s performance on the second subtask for CoI3 test set is higher than CoI’s performance on the second subtask for CoI2. We conjecture that it happens because the previous subtasks can be easier to solve or knowing the subsequent subtask helps improving the performance.
@@
Generalization to Unseen Single Tasks
To assess whether adding compositional instructions helps improve the model’s performance on unseen and difficult single instruction tasks, we tested CoI-tuned models on BIG-Bench Hard (BBH). CoI123-tuned Mistral performs the best (Rouge: 29.39) as shown in Table 4. For Alpaca, the model fine-tuned on CoI12 is also better than the baseline and achieves a higher Rouge score of 28.80. This confirms that having compositional instructions helps the model to understand hard single instruction problems as well.
Generalization to Longer Chains (CoI4,5)
In this experiment, we examine whether our CoI models can generalize to longer chains. We run inference on CoI4 and CoI5 test sets using CoI1, CoI12, and CoI123-tuned Alpaca. [^5] Mistral results are in Table 10 in Appendix Full results. As shown in Figure 6, longer chains ($\sigma=2,3$) in the training set help the model to understand unseen longer chain ($\sigma=4,5$) in the test set as well. Moreover, the performance does not drop as high as CoI1 or the baseline non-fine-tuned models that do not learn the chaining reasoning. We posit that the knowledge of compositional instructions in the training set, even though the length of the chain is shorter than 4 or 5, still helps the model to understand the composed tasks.
Generalization to Downstream Composite Tasks
@@
There has been a notable surge in research focused on fine-tuning LLMs using human instructions. Efrat and Levy (2020) examined LLMs’ ability to follow natural language instructions compared to crowd-workers. Wei et al. (2022a); Sanh et al. (2021) have transformed NLP task descriptions into human-like language instructions and showed that LLMs fine-tuned with those instructions have generalizable capability toward unseen tasks (Chung et al. 2024). Subsequently, many studies have emerged to create new instruction datasets aimed at training models in instruction-tuning paradigm: some instruction datasets are fully written by humans (Wang et al. 2022; Conover et al. 2023), the others are written with the help of LLMs (Honovich et al. 2023; Wang et al. 2023; Taori et al. 2023); some instructions are NLP-specific (Mishra et al. 2022; Wang et al. 2022; Weller et al. 2020), and the others are designed to respond to general-purpose instructions (Ouyang et al. 2022; Wang et al. 2023). These prior studies only work on single instruction datasets, so we construct a new compositional dataset upon Wang et al. (2022)’s Super-NaturalInstruction. Our work is also related to several past works which have leveraged LLMs to generate training data (Schick and Schütze 2021), and some of them specifically use LLMs for generating instruction data (Peng et al. 2023; Shao et al. 2023). Nevertheless, our CoI data generation framework differs from previous works as we use LLMs to determine the composability of individual instructions, and then generate responses for subsequent subtask instructions.